# Explicit Second-Order Min-Max Optimization Methods with Optimal Convergence Guarantee

### v4

@@ Abstract
We propose and analyze several inexact regularized Newton-type methods for finding a global saddle point of convex-concave unconstrained min-max optimization problems. Compared to first-order methods, our understanding of second-order methods for min-max optimization is relatively limited, as obtaining global rates of convergence with second-order information is much more involved. In this paper, we examine how second-order information can be used to speed up extra-gradient methods, even under inexactness. Specifically, we show that the proposed methods generate iterates that remain within a bounded set and that the averaged iterates converge to an $\epsilon$-saddle point within $O(\epsilon^{-2/3})$ iterations in terms of a restricted gap function. This matched the theoretically established lower bound in this context. We also provide a simple routine for solving the subproblem at each iteration, requiring a single Schur decomposition and $O(\log\log(1/\epsilon))$ calls to a linear system solver in a quasi-upper-triangular system. Thus, our method improves the existing line-search-based second-order min-max optimization methods (Monteiro and Svaiter, 2012; Jiang and Mokhtari, 2022) by shaving off an $O(\log\log(1/\epsilon))$ factor in the required number of Schur decompositions. Finally, we present a series of numerical experiments on synthetic and real data that demonstrate the efficiency of the proposed methods.
@@ 1 Introduction
Let $\mathbb{R}^{m}$ and $\mathbb{R}^{n}$ be finite-dimensional Euclidean spaces and assume that the function $f:\mathbb{R}^{m}\times\mathbb{R}^{n}\mapsto\mathbb{R}$ has a bounded and Lipschitz-continuous Hessian. We consider the problem of finding a global saddle point of the following min-max optimization problem:
@@
i.e., a tuple $(\mathbf{x}^{\star},\mathbf{y}^{\star})\in\mathbb{R}^{m}\times\mathbb{R}^{n}$ such that
$f(\mathbf{x}^{\star},\mathbf{y})\leq f(\mathbf{x}^{\star},\mathbf{y}^{\star})\leq f(\mathbf{x},\mathbf{y}^{\star}),\quad\textnormal{for all }\mathbf{x}\in\mathbb{R}^{m},\ \mathbf{y}\in\mathbb{R}^{n}.$
@@
Throughout our paper, we assume that the function $f(\mathbf{x},\mathbf{y})$ is convex in $\mathbf{x}$ for all $\mathbf{y}\in\mathbb{R}^{n}$ and concave in $\mathbf{y}$ for all $\mathbf{x}\in\mathbb{R}^{m}$. This convex-concave setting has been the focus of intense research in optimization, game theory, economics and computer science for several decades now (Von Neumann and Morgenstern, 1953; Dantzig, 1963; Blackwell and Girshick, 1979; Facchinei and Pang, 2007; Ben-Tal et al., 2009), and variants of the problem have recently attracted significant interest in machine learning and data science, with applications in generative adversarial networks (GANs) (Goodfellow et al., 2014; Arjovsky et al., 2017), adversarial learning (Sinha et al., 2018), distributed multi-agent systems (Shamma, 2008), and many other fields; for a wide range of concrete examples, see Facchinei and Pang (2007) and references therein.
@@
Motivated by these applications, several classes of optimization algorithms have been proposed and analyzed for finding a global saddle point of Eq. (1.1) in the convex-concave setting. An important algorithm is the extragradient (EG) method (Korpelevich, 1976; Antipin, 1978; Nemirovski, 2004). The method’s rate of convergence for smooth and strongly-convex-strongly-concave functions and bilinear functions (i.e., when $f(\mathbf{x},\mathbf{y})=\mathbf{x}^{\top}A\mathbf{y}$ for some square, full-rank matrix $A$) was shown to be linear by Korpelevich (1976) and Tseng (1995). Subsequently, Nemirovski (2004) showed that the method enjoys an $O(\epsilon^{-1})$ convergence guarantee for constrained problems with a bounded domain and a convex-concave function $f$. In unbounded domains, Solodov and Svaiter (1999) generalized EG to the hybrid proximal extragradient (HPE) method which provides a framework for analyzing the iteration complexity of several existing methods, including EG and Tseng’s forward-backward splitting (Tseng, 2000), while Monteiro and Svaiter (2010) provided an $O(\epsilon^{-1})$ guarantee for HPE in both bounded and unbounded domains. In addition to EG, there are other methods that can achieve the same convergence guarantees, such as optimistic gradient descent ascent (OGDA) (Popov, 1980) and dual extrapolation (DE) (Nesterov, 2007); for a partial survey, see Hsieh et al. (2019) and the references therein. All these methods are order-optimal first-order methods since they match the lower bound of Ouyang and Xu (2021).
@@
Focusing on convex minimization problems for the moment, significant effort has been devoted to developing first-order methods that are characterized by simplicity of implementation and analytic tractability (Nesterov, 1983). Indeed, it has been recognized that first-order methods are suitable for solving large-scale machine learning problems where low-accuracy solutions may suffice (Sra et al., 2012; Lan, 2020). However, second-order methods are known to enjoy superior convergence properties over their first-order counterparts, in both theory and practice for many other problems: the accelerated cubic regularized Newton method (Nesterov, 2008) and accelerated Newton proximal extra-gradient method (Monteiro and Svaiter, 2013) converge at a global rate of $O(\epsilon^{-1/3})$ and $\tilde{O}(\epsilon^{-2/7})$ respectively, both exceeding the best possible $O(\epsilon^{-1/2})$ bound for first-order methods (Nemirovski and Yudin, 1983). Optimal second-order methods with the rate of $O(\epsilon^{-2/7})$ have been recently proposed by Carmon et al. (2022) and Kovalev and Gasnikov (2022), independently. In addition, first-order methods may perform poorly in ill-conditioned problems and are known to be sensitive to the parameter choices in real-world applications in which second-order methods are observed to be more robust (Pilanci and Wainwright, 2017; Roosta-Khorasani and Mahoney, 2019; Berahas et al., 2020).
@@
In the context of convex-concave min-max problems, two separate issues arise: (i) achieving acceleration with second-order information is less tractable analytically; and (ii) acquiring accurate second-order information is computationally very expensive in general. Aiming to address these issues, a line of recent work has generalized classical first-order methods to their higher-order counterparts (Monteiro and Svaiter, 2012; Bullins and Lai, 2022; Jiang and Mokhtari, 2022), where the best known upper iteration bound is $O(\epsilon^{-2/3}\log\log(1/\epsilon))$ (Jiang and Mokhtari, 2022) with a lower bound of $\Omega(\epsilon^{-2/3})$ (Lin and Jordan, 2024). Closing this gap of $\log\log(1/\epsilon)$ is mainly of theoretical interest but also has important practical implications because all existing methods require a nontrivial implicit search scheme at each iteration, and this can be computationally expensive in practice. [^1] By “implicit,” we mean that the method’s inner-loop subproblem for computing the $k^{\textnormal{th}}$ iterate involves the iterate being updated, leading to an implicit update rule. By contrast, “explicit” means that any inner-loop subproblem for computing the $k^{\textnormal{th}}$ iterate does not involve the new iterate.
In a similar vein, Huang et al. (2022) extended the cubic regularized Newton method (Nesterov and Polyak, 2006) to convex-concave min-max optimization. Their method has two phases and their guarantees require an error bound condition with a parameter $0<\theta\leq 1$ (Huang et al., 2022, Assumption 5.1). In particular, the rate is linear under a Lipschitz-type condition ($\theta=1$) and $O(\epsilon^{-(1-\theta)/\theta^{2}})$ under a Hölder-type condition ($\theta\in(0,1)$). These conditions unfortunately exclude some important problem classes and are hard to verify in general. Another extension of cubic regularized Newton method was provided in Nesterov (2006) and was shown to achieve a global convergence rate of $O(\epsilon^{-1})$ without assuming any error bound condition.
@@
Finally, it is worth mentioning that existing second-order min-max optimization algorithms require the exact second-order information; as a result, given the implicit nature of the inner loop subproblems involved, the methods’ robustness to inexact information cannot be taken for granted. It is thus natural to ask:
@@
Can we develop explicit second-order min-max optimization algorithms that remain order-optimal even with inexact second-order information?
@@
Our paper offers an affirmative answer to the above question. Inspired by recent advances on variational inequalities (VIs) (Lin and Jordan, 2024), we start by presenting a second-order min-max optimization method with a global rate of $O(\epsilon^{-2/3})$. Our convergence analysis here is similar to that of Lin and Jordan (2024) although it is simpler in that it exploits the structure of unconstrained min-max optimization problems. More importantly, our work differs from Lin and Jordan (2024) in that the latter requires the exact second-order information and lacks an explicit complexity analysis for inexact subproblem solving.
@@
Thus, the main contribution of our paper is to relax the requirements of existing work by proposing a class of second-order min-max optimization methods that require only inexact second-order information and inexact subproblem solutions. Moreover, our inexact Jacobian regularity condition allows for the use of randomized sampling for solving finite-sum min-max optimization problems. This yields considerable computational savings since the sample size increases gracefully from a very small sample set. We accordingly prove that the inexact methods achieve a global rate of $O(\epsilon^{-2/3})$ and the subsampled Newton methods achieve the same rate with high probability. Our new subroutine involves solving each subproblem via a single Schur decomposition and $O(\log\log(1/\epsilon))$ calls to a linear system solver in a quasi-upper-triangular system. As such, the total complexity bound of our method is $O((m+n)^{\omega}\epsilon^{-2/3}+(m+n)^{2}\epsilon^{-2/3}\log\log(1/\epsilon))$ where $\omega\approx 2.3728$ is the matrix multiplication constant. Experiments conducted on real and synthetic datasets demonstrate the practical efficiency of this method.
To the best of our knowledge, our method is the first second-order min-max optimization method that does not require exact second-order information. This improves on the existing works (Lin and Jordan, 2024; Adil et al., 2022), which present solutions to convex-concave min-max optimization problems with order-optimal iteration complexity of $\Theta(\epsilon^{-2/3})$, but either require an exact solution of an inner explicit subproblem or lack a characterization of the complexity of solving each subproblem. Indeed, Adil et al. (2022) showed that solving each subproblem requires a single Schur decomposition and $O(\log(1/\epsilon))$ calls to a linear system solver in a quasi-upper-triangular system. Our work is also related to a recent literature on the line-search-based methods (Monteiro and Svaiter, 2012; Bullins and Lai, 2022; Jiang and Mokhtari, 2022; Lin and Jordan, 2023). Differing from these methods, our method is explicit and achieves an order-optimal iteration complexity: specifically, in terms of complexity bound guarantees, our method slightly outperforms the best known line-search-based method (Jiang and Mokhtari, 2022) that achieves the bound of $O((m+n)^{\omega}\epsilon^{-2/3}\log\log(1/\epsilon))$.
@@ 1.1 Related Work
Our work comes amid a surge of interest in optimization algorithms for a large class of emerging min-max optimization problems. For brevity, we will focus on convex-concave settings and leave other settings out of the discussion; see Lin et al. (2020, Section 2) for a more detailed presentation.
@@
Historically, a concrete instantiation of the convex-concave min-max optimization problem is the solution of $\min_{\mathbf{x}\in\Delta^{m}}\max_{\mathbf{y}\in\Delta^{n}}\mathbf{x}^{\top}A\mathbf{y}$ for $A\in\mathbb{R}^{m\times n}$ over the simplices $\Delta^{m}$ and $\Delta^{n}$. Spurred by the von Neumann’s theorem (Neumann, 1928), this problem provided the initial impetus for min-max optimization. Sion (1958) generalized von Neumann’s result from bilinear cases to convex-concave cases and triggered a line of research on algorithms for convex-concave min-max optimization (Korpelevich, 1976; Nemirovski, 2004; Nesterov, 2007; Nedić and Ozdaglar, 2009; Mokhtari et al., 2020b). A notable result is that gradient descent ascent (GDA) with diminishing stepsizes can find an $\epsilon$-global saddle point within $O(\epsilon^{-2})$ iterations if the gradients are bounded over the feasible sets (Nedić and Ozdaglar, 2009).
@@
Recent years have witnessed progress on the analysis of first-order min-max optimization algorithms in bilinear cases and convex-concave cases. In a bilinear case, Daskalakis et al. (2018) proved the convergence of OGDA to a neighborhood of a global saddle point. Liang and Stokes (2019) used a dynamical system approach to establish the linear convergence of OGDA for the special case when the matrix $A$ is square and full rank. Mokhtari et al. (2020a) have revisited proximal point method and proposed an unified framework for achieving the sharpest convergence rates of both EG and OGDA. In the convex-concave case, Nemirovski (2004) demonstrated that EG finds an $\epsilon$-global saddle point within $O(\epsilon^{-1})$ iterations when the feasible sets are convex and compact. The same convergence guarantee was extended to unbounded feasible sets (Monteiro and Svaiter, 2010, 2011) using the HPE method with different optimality criteria. Nesterov (2007) and Tseng (2008) proposed several new algorithms and refined convergence analysis with the same convergence guarantee. Abernethy et al. (2021) presented a Hamiltonian gradient descent method with last-iterate convergence under a “sufficiently bilinear” condition. Focusing on special min-max optimization problems with $f(\mathbf{x},\mathbf{y})=g(\mathbf{x})+\mathbf{x}^{\top}A\mathbf{y}-h(\mathbf{y})$Chambolle and Pock (2011) introduced a primal-dual hybrid gradient method that converges to a global saddle point at the rate of $O(\epsilon^{-1})$ when the functions $g$ and $h$ are simple and convex. Nesterov (2005) proposed a smoothing technique and proved that his algorithm achieves better dependence on problem parameters for convex and compact constraint sets. He and Monteiro (2016) and Kolossoski and Monteiro (2017) proved that such a result still holds true for unbounded feasible sets and non-Euclidean metrics. Moreover, Chen et al. (2014) developed optimal algorithms for solving min-max optimization problems with $f(\mathbf{x},\mathbf{y})=g(\mathbf{x})+\mathbf{x}^{\top}A\mathbf{y}-h(\mathbf{y})$ even when only noisy gradients are available.
@@
Compared to first-order methods, there has been less research on second-order methods for min-max optimization problems with global convergence rate guarantee. In particular, we are aware of two research thrusts (Monteiro and Svaiter, 2012; Bullins and Lai, 2022; Jiang and Mokhtari, 2022; Huang et al., 2022; Adil et al., 2022; Lin and Jordan, 2023, 2024). Our results contribute to this landscape by proposing the first explicit method that achieves the order-optimal iteration complexity of $O(\epsilon^{-2/3})$ and a tight complexity of $O((m+n)^{\omega}\epsilon^{-2/3}+(m+n)^{2}\epsilon^{-2/3}\log\log(1/\epsilon))$. As far as we are aware, the complexity bound guarantees of Algorithms 2 and 3 cannot be realized by other existing second-order min-max optimization methods with exact second-order information requirements.

### v5

# Explicit Second-Order Min-Max Optimization: Practical Algorithms and Complexity Analysis

@@ Abstract
We propose and analyze several inexact regularized Newton-type methods for finding a global saddle point of convex-concave unconstrained min-max optimization problems. Compared to first-order methods, our understanding of second-order methods for min-max optimization is relatively limited, as obtaining global rates of convergence with second-order information can be much more involved. In this paper, we examine how second-order information is used to speed up extra-gradient methods, even under inexactness. In particular, we show that the proposed methods generate iterates that remain within a bounded set and that the averaged iterates converge to an $\epsilon$-saddle point within $O(\epsilon^{-2/3})$ iterations in terms of a restricted gap function. We also provide a simple routine for solving the subproblem at each iteration, requiring a single Schur decomposition and $O(\log\log(1/\epsilon))$ calls to a linear system solver in a quasi-upper-triangular system. Thus, our method improves the existing line-search-based second-order min-max optimization methods (Monteiro and Svaiter, 2012; Bullins and Lai, 2022; Jiang and Mokhtari, 2025) by shaving off an $O(\log\log(1/\epsilon))$ factor in the required number of Schur decompositions. Finally, we conduct experiments on synthetic and real data to demonstrate the efficiency of the proposed methods.
@@ 1 Introduction
Let $\mathbb{R}^{m}$ and $\mathbb{R}^{n}$ be finite-dimensional Euclidean spaces and assume that the function $f:\mathbb{R}^{m}\times\mathbb{R}^{n}\mapsto\mathbb{R}$ has a bounded and Lipschitz-continuous Hessian. We consider the following min-max optimization problem:
@@
and aim to develop methods for finding a global saddle point for this problem; i.e., a tuple $(\mathbf{x}^{\star},\mathbf{y}^{\star})\in\mathbb{R}^{m}\times\mathbb{R}^{n}$ such that
[…]
@@
Throughout this paper, we will assume that the function $f(\mathbf{x},\mathbf{y})$ is convex in $\mathbf{x}$ for all $\mathbf{y}\in\mathbb{R}^{n}$ and concave in $\mathbf{y}$ for all $\mathbf{x}\in\mathbb{R}^{m}$. This convex-concave setting has been the focus of intense research in optimization, game theory, economics and computer science for several decades (Von Neumann and Morgenstern, 1953; Dantzig, 1963; Blackwell and Girshick, 1979; Facchinei and Pang, 2007; Ben-Tal et al., 2009), and variants of this problem class have recently attracted significant interest in machine learning and data science, with applications to generative adversarial networks (GANs) (Goodfellow et al., 2014; Arjovsky et al., 2017), adversarial learning (Sinha et al., 2018), multi-agent systems (Shamma, 2008), and many other fields; see Facchinei and Pang (2007) and references therein for a wide range of concrete problems.
@@
Motivated by these applications, several classes of optimization algorithms have been proposed for finding a global saddle point of Eq. (1.1) in the convex-concave setting. An important algorithm is the extragradient (EG) method (Korpelevich, 1976; Antipin, 1978; Nemirovski, 2004). The method’s rate of convergence for smooth and strongly-convex-strongly-concave functions and bilinear functions (i.e., when $f(\mathbf{x},\mathbf{y})=\mathbf{x}^{\top}A\mathbf{y}$ for some square, full-rank matrix $A$) was shown to be linear by Korpelevich (1976) and Tseng (1995). Subsequently, Nemirovski (2004) showed that the method enjoys an $O(\epsilon^{-1})$ convergence guarantee for constrained problems with a bounded domain and a convex-concave function $f$. In unbounded domains, Solodov and Svaiter (1999) generalized EG to the hybrid proximal extragradient (HPE) method which provides a framework for analyzing the iteration complexity of several existing methods, including EG and Tseng’s forward-backward splitting (Tseng, 2000), while Monteiro and Svaiter (2010) provided an $O(\epsilon^{-1})$ guarantee for HPE in both bounded and unbounded domains. In addition to EG, there are other methods that can achieve the same convergence guarantees, such as optimistic gradient descent ascent (OGDA) (Popov, 1980) and dual extrapolation (DE) (Nesterov, 2007); for a survey, see Hsieh et al. (2019) and the references therein. All these methods are referred to as optimal first-order methods since they have matched the lower bound (Ouyang and Xu, 2021).
@@
Focusing on convex minimization problems for the moment, significant effort has been devoted to developing first-order methods that are characterized by simplicity of implementation and analytic tractability (Nesterov, 1983). In particular, it has been recognized that first-order methods are suitable for solving large-scale machine learning problems in which low-accuracy solutions may suffice (Sra et al., 2012; Lan, 2020). In contrast, second-order methods are known to enjoy superior convergence properties over their first-order counterparts in theory: the accelerated cubic regularized Newton method (Nesterov, 2008) and accelerated Newton proximal extragradient method (Monteiro and Svaiter, 2013) provably converge at a rate of $O(\epsilon^{-1/3})$ and $\tilde{O}(\epsilon^{-2/7})$ respectively, both exceeding the best possible $O(\epsilon^{-1/2})$ bound for first-order methods (Nemirovski and Yudin, 1983). Two optimal second-order methods with the rate of $O(\epsilon^{-2/7})$ have been recently proposed by Carmon et al. (2022) and Kovalev and Gasnikov (2022), independently. In practice, first-order methods may perform poorly in ill-conditioned problems and are known to be sensitive to the parameter choices in real-world applications in which second-order methods are recognized to be more robust and reliable (Pilanci and Wainwright, 2017; Roosta-Khorasani and Mahoney, 2019; Berahas et al., 2020).
@@
In the context of convex-concave min-max problems, two issues arise: (i) achieving acceleration with second-order information is less tractable analytically; and (ii) acquiring accurate second-order information is computationally expensive in general. Aiming to address the first issue, a line of work has generalized practical first-order methods to their higher-order counterparts (Monteiro and Svaiter, 2012; Bullins and Lai, 2022; Jiang and Mokhtari, 2025), where the state-of-the-art iteration bound is $O(\epsilon^{-2/3}\log\log(1/\epsilon))$ (Jiang and Mokhtari, 2025). All these methods require a nontrivial implicit search scheme at every iteration, and this can be prohibitive in practice. [^1] By “implicit,” we mean that the method’s inner-loop subproblem for computing the $k^{\textnormal{th}}$ iterate involves the iterate being updated, leading to an implicit update rule. By contrast, “explicit” means that any inner-loop subproblem for computing the $k^{\textnormal{th}}$ iterate does not involve the new iterate. Another line of work has extended second-order convex optimization methods to their min-max counterparts (Nesterov, 2006; Huang et al., 2022; Chen et al., 2025a). Specifically, the first extension of the cubic regularized Newton method (Nesterov and Polyak, 2006) was provided in Nesterov (2006) and was shown to achieve a global convergence rate of $O(\epsilon^{-1})$Huang et al. (2022) proposed a two-phase extension of cubic regularized Newton method (Nesterov and Polyak, 2006) and provided a theoretical guarantee under the error bound condition with a parameter $0<\theta\leq 1$ (Huang et al., 2022, Assumption 5.1): the global rate is linear under a Lipschitz-type condition ($\theta=1$) and $O(\epsilon^{-(1-\theta)/\theta^{2}})$ under a Hölder-type condition ($\theta\in(0,1)$). These conditions exclude important problem classes and are in general unverifiable. Very recently, Chen et al. (2025a) has shown an improved upper bound of $\tilde{O}(\epsilon^{-4/7})$ iteration bound by generalizing the optimal second-order optimization method (Carmon et al., 2022) to solve convex-concave min-max problems. Despite the appealing theoretical convergence guarantee, this method is a nested-loop algorithm and can thus be complex to implement.
@@
Finally, it is worth mentioning that the existing second-order min-max optimization algorithms require exact second-order information; as a result, given the implicit nature of the inner loop subproblems involved, the methods’ robustness to inexact information cannot be taken for granted. It is thus natural to ask:
@@
Can we develop explicit second-order min-max optimization algorithms that retain the same convergence guarantee even with inexact second-order information?
@@
Our paper offers an affirmative answer to this question. Inspired by recent advances on variational inequalities (VIs) (Lin and Jordan, 2025), we start by presenting a “conceptual” second-order min-max optimization method with an iteration complexity of $O(\epsilon^{-2/3})$. Our convergence analysis closely follows that of Lin and Jordan (2025) although it is simpler in that it leverages the specific structure of unconstrained min-max optimization problems. Although our proof is not new, we include it for completeness.
@@
We then propose a class of second-order min-max optimization methods that require only inexact second-order information and inexact subproblem solutions. Notably, the proposed inexact Jacobian regularity condition allows for using randomized sampling for solving finite-sum min-max problems. This yields considerable computational savings since the sample size increases gracefully from a very small sample set. We prove that our inexact methods can achieve an iteration complexity of $O(\epsilon^{-2/3})$ and the corresponding subsampled variants achieve the same bound with high probability. Our new subroutine for solving each subproblem involves a single Schur decomposition and $O(\log\log(1/\epsilon))$ calls to a linear system solver in a quasi-upper-triangular system. Thus, the total complexity bound is $O((m+n)^{\omega}\epsilon^{-2/3}+(m+n)^{2}\epsilon^{-2/3}\log\log(1/\epsilon))$ where $\omega\approx 2.3728$ is the matrix multiplication constant (Demmel et al., 2007). In addition, we conduct experiments on synthetic and real data to demonstrate the efficiency of the proposed methods.
Our approach builds on existing algorithmic components: (i) the overall scheme is inspired by Lin and Jordan (2025); (ii) the subproblem reformulation is inspired by Adil et al. (2022); (iii) the use of the Newton method and the corresponding iteration complexity bound of $O(\log\log(1/\epsilon))$ extend the analysis of Conn et al. (2000); Cartis et al. (2011); (iv) the subsampling techniques are inspired by Xu et al. (2020). However, we remark that combining these components to yield a theoretically sound and practically efficient algorithm is far from straightforward.
While Adil et al. (2022) showed how to formulate the subproblem as a root-finding problem, a central challenge remains: how to solve such problems efficiently. Techniques such as the generalized conjugate gradient method with Lanczos processes—which are effective in the minimization setting—do not directly extend to the min-max setting. In this context, Huang et al. (2022) proposed an alternative formulation (see Eq. (16) in their paper) and applied a damped Newton method. However, they did not provide theoretical guarantees, even in a local sense, due to the absence of key properties such as the non-singularity of the Jacobian matrix. Even when adopting the root-finding formulation from Adil et al. (2022), identifying a suitable solver is highly nontrivial. At first glance, bisection might seem preferable due to its global convergence guarantee, unlike Newton’s method. One of our major contributions is to rigorously justify the use of the Newton method by extending the analysis of Conn et al. (2000) and Cartis et al. (2011) from the minimization setting to the min-max setting. Moreover, it remains unclear if the existing inexactness conditions from Xu et al. (2020) can be directly combined with our overall scheme while sacrificing neither theoretical iteration complexity nor practical efficiency. For example, the proof of Lemma 4.3 requires separate treatment of $|\Delta\textbf{z}_{k-1}|\geq 1$ and $|\Delta\textbf{z}_{k-1}|<1$, which was unnecessary in previous works either due to stronger assumptions as in Lin and Jordan (2025) or the simpler analysis appropriate for a minimization setting as in Xu et al. (2020).
In summary, our work aims to develop practically effective second-order methods for min-max optimization that maintain the same worst-case iteration complexity bounds. More precisely, we contribute a new class of algorithms that relax some assumptions of prior work by (aiming at practical feasibility) requiring only inexact second-order information and inexact subproblem solutions, while still achieving same theoretical performance. This improves on existing works (Lin and Jordan, 2025; Adil et al., 2022), which either require an exact solution of an inner subproblem or lack a complexity characterization of solving each subproblem. In terms of complexity guarantees, our method outperforms the best line-search-based method (Jiang and Mokhtari, 2025) that achieves the bound of $O((m+n)^{\omega}\epsilon^{-2/3}\log\log(1/\epsilon))$. Subsequent to our work, new strategies were designed for improving second-order min-max optimization methods (Alves et al., 2024; Jiang et al., 2024; Chen et al., 2025b), thereby matching and even surpassing our results. While we believe these strategies could be used in our method to yield an improved guarantee, we consider it beyond the scope of this work and leave it for future research.
@@ 1.1 Related Work
Our work comes amid a surge of interest in optimization algorithms for a large class of emerging min-max optimization problems. For brevity, we focus on convex-concave settings and will leave other settings aside; see Section 2 in Lin et al. (2020) for a more thorough presentation of algorithms and complexity results.
@@
Historically, a concrete instantiation of the convex-concave min-max optimization problem is the solution of $\min_{\mathbf{x}\in\Delta^{m}}\max_{\mathbf{y}\in\Delta^{n}}\mathbf{x}^{\top}A\mathbf{y}$ for $A\in\mathbb{R}^{m\times n}$ over the simplices $\Delta^{m}$ and $\Delta^{n}$. Spurred by von Neumann’s theorem (Neumann, 1928), this problem provided the initial impetus for min-max optimization. Sion (1958) generalized von Neumann’s result from bilinear cases to convex-concave cases and triggered a line of research on algorithms for convex-concave min-max optimization (Korpelevich, 1976; Nemirovski, 2004; Nesterov, 2007; Nedić and Ozdaglar, 2009; Mokhtari et al., 2020b). A notable result is that gradient descent ascent (GDA) with diminishing stepsizes can find an $\epsilon$-global saddle point within $O(\epsilon^{-2})$ iterations if the gradients are bounded over the feasible sets (Nedić and Ozdaglar, 2009).
@@
Recent years have witnessed much progress on the design and analysis of first-order min-max optimization algorithms in bilinear cases and convex-concave cases. In the bilinear case, Daskalakis et al. (2018) proved the convergence of OGDA to a neighborhood of a global saddle point. Liang and Stokes (2019) used a dynamical system approach to establish the linear convergence of OGDA for the special case when the matrix $A$ is square and full rank. Mokhtari et al. (2020a) revisited a proximal point algorithmic framework and proposed an unified analysis for achieving the sharpest convergence rates of both EG and OGDA. In the convex-concave case, Nemirovski (2004) demonstrated that EG finds an $\epsilon$-global saddle point within $O(\epsilon^{-1})$ iterations when the feasible sets are convex and bounded. The same convergence guarantee was extended to unbounded sets (Monteiro and Svaiter, 2010, 2011) using the HPE method with different criteria. In a similar vein, Nesterov (2007) and Tseng (2008) proposed new algorithms and refined convergence analysis with the same convergence guarantee. Abernethy et al. (2021) presented a Hamiltonian gradient descent method with last-iterate convergence under a “sufficiently bilinear” condition. Focusing on a special class of min-max optimization problems with separable structure: $f(\mathbf{x},\mathbf{y})=g(\mathbf{x})+\mathbf{x}^{\top}A\mathbf{y}-h(\mathbf{y})$Chambolle and Pock (2011) introduced a primal-dual hybrid gradient method that converges to a global saddle point at the rate of $O(\epsilon^{-1})$ when the functions $g$ and $h$ are simple and convex. Nesterov (2005) introduced a smoothing technique and proved that the resulting algorithm achieved better dependence on some problem parameters for convex and bounded feasible sets. He and Monteiro (2016) and Kolossoski and Monteiro (2017) proved that such a result holds for unbounded sets and non-Euclidean metrics. In addition, Chen et al. (2014) proposed a new class of stochastic algorithms for solving separable min-max optimization problems when only noisy gradients are available and proved the optimal convergence guarantee.
There are also several papers on finite-sum min-max optimization and variance reduction (Balamurugan and Bach, 2016; Chavdarova et al., 2019; Carmon et al., 2019; Yang et al., 2020; Luo et al., 2021; Alacaoglu and Malitsky, 2022). In the strongly-convex-strongly-concave cases, Balamurugan and Bach (2016) studied stochastic variance-reduced variants of forward-backward methods and proved linear convergence. Chavdarova et al. (2019) combined EG with variance reduction and proved new convergence guarantee which was however less favorable than Balamurugan and Bach (2016). Recently, Luo et al. (2021) have considered unbalanced strongly-convex-strongly-concave cases and proposed variance-reduced first-order methods with an improved complexity bound. In the bilinear case, Carmon et al. (2019) proposed a randomized mirror-prox method with an improved complexity guarantee. However, extending their results to general convex-concave cases requires additional restrictive assumptions. Alacaoglu and Malitsky (2022) addressed this issue by leveraging a new variance reduction technique and their methods either match or improve the best-known complexities for solving convex-concave finite-sum min-max problems. Beyond the convex-concave cases, Yang et al. (2020) provided various variance-reduced first-order methods for solving structured nonconvex-concave min-max optimization problems.
@@
Compared to first-order methods, there has been less research on second-order methods for min-max optimization problems with global convergence rate guarantee. In this context, we are aware of two research thrusts (Monteiro and Svaiter, 2012; Bullins and Lai, 2022; Huang et al., 2022; Adil et al., 2022; Lin and Jordan, 2023, 2025; Jiang and Mokhtari, 2025; Chen et al., 2025a). Our results contribute to this landscape by proposing the first explicit method that achieves the iteration complexity of $O(\epsilon^{-2/3})$ as well as a new complexity bound of $O((m+n)^{\omega}\epsilon^{-2/3}+(m+n)^{2}\epsilon^{-2/3}\log\log(1/\epsilon))$. As far as we know, the above complexity bound guarantees cannot be realized by other existing second-order min-max optimization methods with exact second-order information requirements. Subsequent to our work, new strategies were designed for improving second-order min-max optimization methods (Alves et al., 2024; Jiang et al., 2024; Chen et al., 2025b), thereby matching and even surpassing our results.